What do you mean by acceleration odue to gravity state the universal law pf gravitation....

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Answer:

The 9.8 m/s^2 is theaccelerationof an objectdue to gravityat sea level on earth. You get this value from theLawofUniversal Gravitation. Force = m*a = G(M*m)/r^2. Here you use the radius of the earth for r, the distance to sea level from the center of the earth, and M is the mass of the earth.

Thanks